This is a percents system of equations word problem. Please help me solve it.

A chemistry student needs 40 milliliters (mL) of a 14% acid solution. She had two acid solutions, A and B, to mix together to form the 40 mL acid solution. Acid solution A is 10% and acid solution B is 20% acid. How much of each solution should be used?
I have created one equation: x + y = 40
The second equation I have written is not finished:
.1x + .2y = ????
How do I finish the second equation? And why?

Answer by ankor@dixie-net.com(22740)

A chemistry student needs 40 milliliters (mL) of a 14% acid solution.
She had two acid solutions, A and B, to mix together to form the 40 mL acid solution.
Acid solution A is 10% and acid solution B is 20% acid.
How much of each solution should be used?
I have created one equation: x + y = 40
The second equation I have written is not finished:
It should be:
.1x + .2y = .14(40)
.1x + .2y = 5.6
:
Using the first equation
x + y = 40
x = (40-y)
:
In the 2nd equation, replace x with (40-y)
.1(40-y) + .2y = 5.6
4 - .1y + .2y = 5.6
.1y = 5.6 - 4
.1y = 1.6
y = 16 liters of 20% solution
:
I'm sure you can find x.